OpenStudy (anonymous):

(x+6)(x-6) multiply and simplify

OpenStudy (anonymous):

(x+6)(x-6) (a+b)(a-b) = a^2-b^2 x^2-6^2= x^2-36 ans

OpenStudy (anonymous):

x^2-36=0